OpenStudy (anonymous):

solve the equation . state the solution set and check yout answer when appropriate 0.12(a + 2) / 3 = 0.114

OpenStudy (anonymous):

how do you solve it out?

OpenStudy (mathmale):

while not strictly necessary, multiplying both sides of this equation by 1000 will eliminate the (decimal) fractions. Next, I'd suggest you multiply both sides of the resulting equation by 3, to remove that denominator of 3. Would you please do these two things now. Type out your answers.

OpenStudy (anonymous):

120 (a + 2) / 3 = 114 then 360 (a + 2) =342 @mathmale

OpenStudy (mathmale):

120 (a + 2) / 3 = 114 is fine. Multiply both sides by 3, noting that the 3 in the numerator on the left side will cancel the 3 in the denominator on the left. 360 is not correct. (Why?)

OpenStudy (anonymous):

oh I multiplied the 120 by the instead of canceling it out

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Yes. So, what is the correct result after you've multiplied both sides of the equation by 3?

OpenStudy (anonymous):

120 (a + 2) = 342 @mathmale

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Looks good. So much simpler without those fractions! since this is an equation, you can proceed to solve it for the value of 'a'. Know how to do that? First, you might want to simplify this equation by finding the largest common factor of 120 and 342 and dividing both by that LCF.

OpenStudy (mathmale):

Hint: is 3 a factor of both 120 and 342? Is 3 the GCF (greatest common factor)?

OpenStudy (anonymous):

i did it and i got a=220

OpenStudy (mathmale):

You could now check your answer (a=220) by substituting it back into the original equation. Does doing this make the original equation true? if so, you know a=220 is the solution.

OpenStudy (anonymous):

i checked it and got 8.88 did i do something wrong

OpenStudy (mathmale):

\[0.12(a + 2) / 3 = 0.114\]Multiply both sides by 100 to obtain\[120(a+2)/3=114\]

OpenStudy (mathmale):

Next, \[120(a+2)=342\]

OpenStudy (mathmale):

Since we're looking to find the value of 'a,' please just divide both sides of this equation by 120. What do you obtain by doing that?

OpenStudy (anonymous):

2.85?

OpenStudy (mathmale):

If you do that, you end up with a+2=2.85, right? Try subtracting 2 from both sides of this equation. What do you get?

OpenStudy (anonymous):

a=.85

OpenStudy (anonymous):

thank you

OpenStudy (mathmale):

Nice work. Hope to work with you again soon!